You made a square card to a friend. The card did not fit in the envelope so you had to trim the card. You trimmed 4 inches from the length and 5 inches from the width. The area of the resulting card is 20 square inches. What were the original dimensions of the card?
Answer by richwmiller(17219)

(x-4)*(x-5)=20
x^2-9x+20=20
x^2-9x=0
x^2=9x
x=9 inch square card